# Quantum-Enhanced LLM Efficient Fine Tuning

### v1

@@ III-B Evaluation Metrics
In this experiment, Cross-Entropy Loss[23] is adopted as the evaluation metric for the model. Cross-entropy measures the difference between the predicted probability distribution of the model and the true label distribution, effectively reflecting the optimization level of classification tasks. A smaller value indicates that the model’s predictions are closer to the true distribution. For a dataset containing N samples, the cross-entropy loss is calculated as:
@@
For text generation evaluation, we utilize BLEU-4[24] and ROUGE[25] scores. BLEU-4 calculates 4-gram precision with a brevity penalty BP to penalize short outputs:
[…]
$\text{BP}=\begin{cases}1&\text{if }c>r\\ e^{(1-r/c)}&\text{if }c\leq r\end{cases},$
$\text{Precision}_{n}=\frac{\sum\text{Count}_{\text{clip}}(\text{n-gram})}{\sum\text{Count}_{\text{generated}}(\text{n-gram})},$
  • BLEU-4: Bilingual Evaluation Understudy score for 4-gram precision,
  • BP: Brevity Penalty to penalize short translations,
  • $w_{n}$: Uniform weights with $w_{n}=\frac{1}{4}$,
  • $\text{Precision}_{n}$: Modified n-gram precision for $n=1$ to $4$,
  • $c$: Total length of candidate translations,
  • $r$: Effective reference corpus length,
[…]
The recall metric ROUGE-N quantifies the proportion of n-grams in the reference(s) text that are successfully reproduced in the generated text. Formally, it is computed as:
[…]
  • R: the set of reference text(s),
[…]
The ROUGE-L metric evaluates text generation quality by measuring the alignment between a generated text and reference text(s) using the Longest Common Subsequence (LCS). Unlike n-gram-based metrics, ROUGE-L emphasizes semantic coherence and structural continuity by identifying the longest sequence of tokens that appear in both texts in the same order, though not necessarily contiguously. The metric is computed as a balanced F1-score derived from LCS-based recall ($R_{\text{LCS}}$) and precision ($P_{\text{LCS}}$):
[…]
$R_{\text{LCS}}=\frac{\text{Length of LCS(Ref,Gen)}}{\text{Total tokens in reference text(s)}},$
$P_{\text{LCS}}=\frac{\text{Length of LCS(Ref,Gen)}}{\text{Total tokens in generated text}},$
  • $R_{\text{LCS}}$: Recall of the Longest Common Subsequence,
  • $P_{\text{LCS}}$: Precision of the Longest Common Subsequence,
  • $\beta$: Controls the relative weight of recall ($R_{\text{LCS}}$) and precision ($P_{\text{LCS}}$). Typically, $\beta$=1 to prioritize harmonic mean (F1).
III-C Based Model
The baseline model for comparison is the LoRA algorithm with a rank of 4, fine-tuned based on the DeepSeek-R1-Distill-Qwen-7B (https://modelscope.cn/models/deepseek-ai/DeepSeek-R1-Distill-Qwen-7B) and Qwen2-7B-Instruct models.
III-D Results
@@
This section primarily aims to verify whether the efficient fine-tuning algorithm described for the QWTHN can improve the performance of fine-tuned models while significantly reducing the number of parameters. This evaluation is based on the assumption that the QWTHN can enhance the expressive power of classical neural networks in low-rank spaces. The experiments were conducted using PyTorch[26] and PyVQNet[27, 28] with quantum circuit simulations. In this stage, noise effects in the quantum system are ignored, and the Cross-Entropy Loss is adopted as the optimization objective for classification tasks. As shown in Fig 3, the convergence curve of QWTHN exhibits a steeper descent rate during early training stages and achieves a lower stable loss value in training phase.
@@
Figure 3: Training loss comparison of QWTHN, QuanTA, and LoRA across varying sample sizes (N) and datasets. Subplots (a–c) illustrate results on the CPsyCounD dataset with N={300,600,1000}, while subplots (d–f) show corresponding analyses on the R1-Distill-SFT dataset. QWTHN consistently achieves lower training losses compared to baseline methods, demonstrating enhanced convergence efficiency under diverse experimental conditions.
@@
The datasets are extracted from CPsyCounD and R1-Distill-SFT with 300, 600, and 1000 samples, respectively. The maximum sequence length max_seq_len for both training and prediction is set to 1024, with 10% of the data used as the validation set and the remaining 90% as the training set. For the Qwen2-7B-Instruct test set, the last 50 samples from CPsyCounD (with no overlap with the previous data) are used as the test set.
Since DeepSeek-R1-Distill-Qwen-7B excels in code and mathematical capabilities, using metrics like BLEU-4 to compare the performance of QWTHN LoRA and MLP LoRA is not appropriate. Instead, we evaluate the test set performance using the following metrics:
@@
Strict Accuracy (SA): Defined as whether the output contains the correct answer in the answer segment. Accuracy: Unlike SA, this metric considers the answer correct if the correct answer appears anywhere in the reasoning chain. Chain-of-Thought Completeness (CTC): Evaluates whether the response includes a complete reasoning chain, assessing the model’s ability to apply reasoning chains within limited lengths. Answer Completeness (AC): Measures the completeness of the answer itself.
@@
As demonstrated in Table I, the proposed QWTHN achieves significantly higher training efficiency than classical LoRA under identical fine-tuning conditions for LLMs. When evaluated on chain-of-thought (CoT) reasoning tasks, QWTHN exhibits statistically significant improvements in domain-specific metrics—including Contextual Task Consistency (CTC), Answer Coherence (AC), Accuracy, and Strict Accuracy—while reducing trainable parameters by over 70% compared to LoRA.
@@
Further validation across benchmark datasets (CPsyCounD, R1-Distill-SFT, and Chinese-DeepSeek-R1-Distill-data-110k) confirms the robustness of QWTHN. As summarized in Table II, the framework demonstrates consistent gains in text generation quality, with absolute increases of 1–3 points in BLEU-4, ROUGE-1, ROUGE-2, and ROUGE-L metrics. These results underscore QWTHN’s ability to reconcile parameter efficiency with enhanced model performance, establishing it as a viable alternative to conventional low-rank adaptation paradigms.
Table I: Cross-Architecture Performance Comparison by Dataset Scale
@@
Model Small (300) Medium (600) Full (1000)
@@
Model Small (300) Medium (600) Full (1000)
@@
MLP 0.179 0.182 0.172 0.158 0.169 0.160
quanTA 0.167 0.173 0.164 0.159 0.163 0.159
QWTHN 0.155 0.174 0.158 0.166 0.162 0.164
Note. Parameter counts: MLP = 1.26M (100%), quanTA = 1.37M (108.7%), QWTHN = 0.30M (23.7%).
Table II: Performance Metrics for CPsyCounD, R1-Distill-SFT and Chinese-DeepSeek-R1-Distill-data-110k
@@
MLP 13.459 34.540 13.109 30.627
quanTA 8.650 29.463 9.780 25.539
QWTHN 14.593 37.345 14.023 32.32
R1-Distill-SFT
Model BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
MLP 59.635 64.195 42.627 43.326
quanTA 59.096 63.512 41.992 42.227
QWTHN 61.615 66.087 45.294 45.915
Chinese-DeepSeek-R1-Distill-data-110k
@@
MLP 78% 68% 82% 68%
quanTA 84% 74% 82% 68%
QWTHN 96% 94% 86% 82%
III-E Quantum Running
We have constructed a variational quantum circuit module incorporating the Origin Quantum Cloud (QCloudService) in our model architecture. This module utilizes the quantum chip backend codenamed "origin_wukong" to transform user-defined variational quantum circuits (including input data and trainable parameters) into quantum circuit intermediate representations (originIR) compatible with the QPanda framework[29]. Given that Origin Quantum chips exclusively support Z-axis observable measurements, basis transformation operations are implemented to convert user-defined non-Z observables into equivalent Z-axis measurements. Specifically, each Pauli operator measurement generates a distinct quantum circuit.
@@ III-E Quantum Running
Finally, measurement results are reformatted into machine learning-compatible tensors through concatenation and reshaping operations. These processed outputs can be directly utilized by subsequent neural network modules, establishing seamless integration between quantum computations and classical machine learning pipelines.
III-F Discussion
The parameter count of the Quantum Tensor Hybrid Network (QWTHN), which combines quantum tensor networks and quantum neural networks, is lower than that of classical LoRA, while the training loss is significantly reduced. This leads to improved parameter efficiency. The Matrix Product Operator (MPO) decomposition introduces the concept of rank, effectively decomposing high-dimensional linear transformations into a series of low-dimensional matrix products. This decomposition significantly reduces the number of model parameters, especially when handling high-dimensional inputs and outputs.
Traditional fully connected layers flatten the input into vectors, losing the dimensional structure of the original data. In contrast, MPO neural networks preserve this structural information to some extent because they operate on the dimensions of the original tensor. This structured approach may enable MPO to learn and represent multidimensional relationships and patterns in the data more effectively, such as correlations between dimensions and hierarchical structures.
By leveraging the classical output of the QNN to perform element-wise multiplication with the output of the classical network, the expressive power of classical neural networks in low-rank spaces is enhanced. Quantum neural networks process input data using quantum properties (e.g., superposition, entanglement, and parallel computation), potentially generating high-dimensional or nonlinear features. Even though the final output of the QNN is classical data (e.g., probability distributions after measurement), the underlying quantum state evolution process can still extract information that is difficult to obtain using classical methods.
Currently, the scale of dataset experiments is limited by the size of quantum circuits. Although it can be seen that training loss on a larger dataset of 1000 samples still performs better. Future work will need to focus on developing fine-tuning strategies for larger datasets and even exploring large-scale quantum pre-training schemes. The current integration of MPO (Matrix Product Operator) and VQC (Variational Quantum Circuit) involves hard connections between input and output interactions. However, theoretically, all weight matrices could be replaced with variational quantum circuits, which would enable the model to capture higher degrees of correlations far beyond what classical weight matrices can represent. Weight matrix W can be decomposed into an MPO with a given bond dimension, and this MPO can be further disentangled using quantum circuits, with the remaining operators also expressed in MPO format [30]. This is one feasible technical direction, and we will explore more potential implementations of large-scale quantum models.
@@ IV Conclusion
This paper implements a quantum-weighted tensor hybrid fine-tuning algorithm for large models, leveraging quantum computing principles to significantly reduce parameter redundancy and enhance parameter efficiency. It validates the potential of quantum computing to alleviate the computational bottleneck of large models, achieving a novel integration of quantum computing and large-scale models. At the current stage of quantum computer development, this work not only demonstrates the efficient adaptation of quantum resources to billion-parameter models for lightweight optimization but also validates a practical pathway for quantum hardware driven by large model tasks. This lays the first engineerable technical foundation for future quantum-enhanced Artificial General Intelligence (AGI) systems.

### v2

@@ III-B Evaluation Metrics
In this experiment, cross-entropy loss[23] is adopted as the training and validation loss function for the model. Cross-entropy measures the difference between the predicted probability distribution of the model and the true label distribution, effectively reflecting the optimization level of classification tasks. A smaller value indicates that the model’s predictions are closer to the true distribution. For a dataset containing $N$ samples, the cross-entropy loss is calculated as:
@@
Perplexity (PPL) is a common evaluation metric in natural language processing used to measure how well a language model predicts a text sequence. It reflects the model’s uncertainty when assigning probabilities to words in the test data. Lower perplexity indicates better performance. PPL is defined as:
$\text{PPL}=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P(w_{i}|w_{<i})\right),$
where $N$ denotes the total number of words/tokens in the test data and $P$ is the probability assigned by the model to the $i$-th word given prior context. A lower PPL means that the model assigns higher probabilities to the correct words, showing better alignment with the true data distribution.
For the evaluation of text generation, we use BLEU-4 [24] and ROUGE [25]. BLEU-4 assesses lexical precision through 4-gram matching between generated and reference texts, while ROUGE measures content coverage and semantic coherence using recall-oriented n-gram and sequence alignment metrics. This combination provides complementary perspectives on the quality of the generation.
III-C Results
Benchmark of various fine-tuning methods using DeepSeek-R1-Distill-Qwen-7B [26] and Qwen2-7B-Instruct [27] models as the base model.The ranks of LoRA [1] and QTHA are set to 4, and the decomposition factor of QuanTA [5] is set to 5. QTHA, LoRA and QuanTA are applied to the linear projection layers q_proj and v_proj.
@@ III-C Results
This section primarily aims to verify whether the efficient fine-tuning algorithm described for the QTHA can improve the performance of fine-tuned models while significantly reducing the number of parameters. The experiments were conducted using PyTorch [28] and PyVQNet [29, 30] with quantum circuit simulations. As shown in Fig. 2, the convergence curve of QTHA exhibits a steeper descent rate during early training stages and achieves a lower stable loss value in training phase.
@@
Figure 2: Training loss comparison of QTHA, LoRA and QuanTA across varying sample sizes (N) and datasets. Subplots (a–b) illustrate results on the CPsyCounD dataset, while subplots (c–d) and (e–f) show corresponding analyses on the R1-Distill-SFT and CH-R1-Math dataset. QTHA consistently achieves lower training losses compared to baseline methods, demonstrating enhanced convergence efficiency under diverse experimental conditions.
@@
The datasets are extracted from CPsyCounD, R1-Distill-SFT, and CH-R1-Math, with 300, 600, 1000, 3000 respectively. The maximum sequence length max_seq_len for both training and prediction is set to 1024, with 10% of the data used as the validation set and the remaining 90% as the training set. For the test set, 30 to 100 samples from the aforementioned three datasets are extracted as the test set.
DeepSeek-R1-Distill-Qwen-7B exhibits excellent performance in code and mathematical reasoning tasks. we use the following metrics to perform a small-scale evaluation of the performance of the test set:
  • Strict Accuracy (SA): Defined as whether the output contains the correct answer in the answer segment.
  • Accuracy: Unlike SA, this metric considers the answer correct if the correct answer appears anywhere in the reasoning chain.
@@
  • Chain-of-Thought Completeness (CTC): Evaluates whether the response includes a complete reasoning chain, assessing the model’s ability to apply reasoning chains within limited lengths.
  • Answer Completeness (AC): Measures the completeness of the answer itself.
@@
As demonstrated in Table I, the proposed QTHA achieves significantly higher training efficiency than classical LoRA under identical fine-tuning conditions for LLMs. When evaluated on chain-of-thought (CoT) reasoning tasks, QTHA exhibits statistically significant improvements in domain-specific metrics—including Contextual Task Consistency (CTC), Answer Coherence (AC), Accuracy, and Strict Accuracy—while reducing trainable parameters by over 76% compared to LoRA.
@@
Further testing on benchmark datasets (CPsyCounD, R1-Distill-SFT, and CH-R1-Math) confirms the robustness of QTHA. As summarized in Table II and III, the framework demonstrates consistent improvements in text generation quality, with notable enhancements observed in the BLEU-4, ROUGE-1, ROUGE-2, ROUGE-L, PPL metrics.These results underscore QTHA’s ability to reconcile parameter efficiency with enhanced model performance, establishing it as a viable alternative to conventional low-rank adaptation paradigms.
Table I: Cross-Architecture Performance Comparison by Dataset Scale.
@@
Dataset Scale LoRA QuanTA QTHA
@@
Dataset Scale LoRA QuanTA QTHA
@@
300 0.179 0.182 0.173 0.178 0.160 0.169
600 0.172 0.162 0.170 0.163 0.159 0.158
1000 0.169 0.160 0.168 0.163 0.160 0.159
3000 0.229 0.213 0.220 0.208 0.211 0.202
CH-R1-Math (Training/Validation Loss)
Dataset Scale LoRA QuanTA QTHA
Train Val Train Val Train Val
300 0.880 0.876 0.855 0.853 0.804 0.804
600 0.829 0.788 0.821 0.791 0.776 0.767
1000 0.800 0.765 0.804 0.781 0.773 0.763
3000 0.852 0.819 0.830 0.803 0.789 0.762
Note. Parameter counts: LoRA = 1.26M (100%), QuanTA = 0.73M (57.5%), QTHA = 0.30M (23.7%), additionally, the number of parameters trained based on R1-Distill-SFT is 0.57M(45.2%).
Table II: Performance Metrics for CPsyCounD, R1-Distill-SFT and CH-R1-Math. Here, “Simulator” denotes the results of QTHA produced by a quantum circuit simulator; and “Wukong” produced by the superconducting quantum computer “Origin Wukong”.
@@
LoRA 14.823 39.560 17.216 34.239
QuanTA 13.368 37.860 15.257 33.705
Simulator 16.043 40.638 17.614 35.892
Wukong 18.056 42.668 20.322 37.172
CH-R1-Math
@@
LoRA 90 % 83 % 97 % 83 %
QuanTA 97 % 93 % 97 % 93 %
Simulator 100 % 100 % 93 % 93 %
Wukong 100 % 100 % 100 % 100 %
Table III: PPL for CPsyCounD, R1-Distill-SFT and CH-R1-Math. “Simulator” or “Wukong” means QTHA executed on a quantum circuit simulator or “Origin Wukong”, respectively.
Model PPL(CPsyCounD) PPL(CH-R1-Math) PPL(R1-Distill-SFT)
LoRA 5.5785 2.5235 1.3355
QuanTA 5.0758 2.4675 1.3306
Simulator 4.8350 2.3556 1.3241
Wukong 4.8325 2.3550 1.3237
III-D Implementation on quantum hardware
QTHA demonstrates excellent performance on noisy quantum computers, while the output of LLMs is inherently a probability distribution matrix, its impact on the final result is always constrained within a specific threshold range, naturally exhibiting stochastic characteristics. The inherent random noise in quantum systems (such as decoherence and gate operation errors) may enhance reasoning robustness through the following mechanisms: in the probability matrix output by LLMs, only when the probability of a key token exceeds a preset threshold will it significantly influence decision-making. This implies that minor perturbations introduced by noise may be filtered out by these thresholds, thereby maintaining result stability. The probabilistic nature of quantum systems (e.g., superposition state collapse) and the probabilistic output of LLMs share mathematical similarities, and noise may implicitly calibrate the probability distribution to improve reasoning consistency.
We have constructed a variational quantum circuit module incorporating the Origin Quantum Cloud (QCloudService) in our model architecture. This module uses a quantum cloud service with a superconducting quantum computer backend named “Origin Wukong”, transforming user-defined variational quantum circuits (including input data and trainable parameters) into quantum circuit intermediate representations (OriginIR) [31]. Given that the backend exclusively support Z-axis observable measurements, basis transformation operations are implemented to convert user-defined non-Z observables into equivalent Z-axis measurements. Specifically, each Pauli operator measurement generates a distinct quantum circuit.
@@ III-D Implementation on quantum hardware
Finally, measurement results are reformatted into machine learning-compatible tensors through concatenation and reshaping operations. These processed outputs can be directly processed by subsequent neural network modules, establishing seamless integration between quantum computations and classical machine learning pipelines.
@@ IV Conclusion
This paper introduces QTHA, a quantum tensor hybrid adaptation framework for fine-tuning LLMs, achieving significant breakthroughs in parameter reduction and performance enhancement by integrating quantum tensor networks with quantum neural networks. Leveraging MPO decomposition, the framework transforms high-dimensional linear transformations into low-dimensional matrix product sequences via rank constraints, drastically reducing parameter counts while preserving tensor structures to effectively capture multi-dimensional data correlations. Additionally, the QNN generates high-dimensional nonlinear features unattainable by classical methods, and its fusion with the MPO network significantly enhances representation power in low-rank spaces. Experimental results demonstrate superior training loss reduction, even with limited datasets of 3,000 samples, highlighting quantum computing’s potential to address computational bottlenecks in LLMs. Looking forward, the research will explore large-scale quantum pre-training schemes and fully quantum weight matrix reconstruction using variational quantum circuits to capture higher-order correlations. This work not only achieves lightweight optimization for billion-parameter models but also establishes a practical paradigm where LLM-driven tasks advance quantum hardware development. It lays the foundation for future quantum-enhanced Artificial General Intelligence systems, marking a key step toward scalable and engineerable quantum AI solutions.